OpenStudy (anonymous):

Find all solutions in the interval [0,2pi). sec^2x-2=tan^2x

OpenStudy (anonymous):

sec^2x or sec(2x) please be clear

OpenStudy (anonymous):

sec^2x

OpenStudy (anonymous):

sec^2x-2=tan^2x tan^2 x -1=tan^2 x -1=0 no solution

OpenStudy (anonymous):

Thank you so much!